MOLECULAR ION SPECTROSCOPY OF BACL^{+}

\underline{STEVEN J. SCHOWALTER}, KUANG CHEN, WADE G. RELLERGERT, SCOTT T. SULLIVAN, AND ERIC R. HUDSON, Department of Physics and Astronomy, University of California, Los Angeles, California 90095, USA.

We discuss our efforts to perform high-resolution spectroscopy of the BaCl^{+} ion, an exciting candidate for ultracold molecular ion studies. This work details our search for a predicted predissociation channel between the first-excited B^{1}Σ and A^{1}Π states. It is expected that the rovibrational resolution afforded by predissociation spectroscopy will allow us to efficiently measure molecular-ion rovibrational temperatures. This is a crucial step in confirming our method to produce ultracold molecular ions via sympathetic collisions with a ^{40}Ca MOT. To observe the predissociation of trapped BaCl^{+}, we detect slight increases in fragment Ba^{+} with a novel time-of-flight device using radial extraction from a linear quadrupole trap.